# Training Small Reasoning LLMs with Cognitive Preference Alignment

### v1

@@ 3.2.3 Verifier
Finally, we leverage a LLM agent $\pi_{\text{large}}$ to validate the correctness and the rewritten reasoning processes, in order to preserve the high quality of the dataset. It predicts whether $\pi_{\text{base}}$ can derive the correct answer $y$ from the rewritten thoughts for $*r_{\text{easy*}}$ or $*r_{\text{hard*}}$ produced by the Rethinker.
After the verification, incorrect cases are sent back to the Rethinker to be continuously rewritten until they pass verification. In the implementation, a maximum of three iterations are allowed; cases that fail to pass verification after three iterations are discarded. As for the correct cases, we invoke the Critic to make the judgment again (please refer to Figure 2 for the algorithmic flow). The final SFT dataset is composed of verified data: $*\mathcal{D}_{\text{SFT*}}=\mathcal{D}_{\text{med}}\cup\mathcal{D}_{\text{easy*}}\cup\mathcal{D}_{\text{hard*}}$, where $*\mathcal{D}_{\text{easy*}}$ and $*\mathcal{D}_{\text{hard*}}$ represent the rewritten versions of $\mathcal{D}_{\text{easy}}$ and $\mathcal{D}_{\text{hard}}$ that have passed verification and have been re-rated as medium by the Critic, respectively. $*\mathcal{D}_{\text{SFT*}}$ serves as the training set for the first stage of CRV distillation training via SFT.
@@ 3.3 Cognitive Preference Optimization
The CogPO algorithm aligns the thought processes of smaller LLMs with their inherent cognitive capacities, following the direct SFT distillation of CoT results refined with the CRV framework.
@@ 3.3.1 Preliminaries
Briefly speaking, CogPO is extended from DPO Rafailov et al. (2023) and its variants. Let $y_{w}$ and $y_{l}$ be the chosen and rejected responses for an instruction $x$ (not restricted to reasoning problems addressed in this work), respectively. We further denote $\pi_{\theta}$ as the model to be optimized after SFT and $\pi_{\text{ref}}$ as the reference model. DPO seeks to maximize the following margin:
@@
The settings of $\beta$ are critical to the performance of DPO. $\beta$-DPO Wu et al. (2024a) further adjusts $\beta$ according to $M_{\beta}(x,y_{w},y_{l})$ either at the instance level or batch level, allowing the model to adapt $\beta$ based on the reward differential of the input data.
@@ 3.3.2 Algorithmic Description
As noted, DPO and $\beta$-DPO do not require any prior knowledge of how the model learns the user’s preferences. We suggest that this type of prior knowledge is critical for training better smaller reasoning models, as the cognitive trajectories of large and small models often differ Li et al. (2022); Zhang et al. (2024); Hu et al. (2024), which may not be directly reflected in the reward differential. We propose CogPO to align reasoning preferences by encoding more prior knowledge, continuously training on a series of mini-tasks.
Recall that it is probable for the Rethinker in CRV to generate incorrect reasoning processes when asked to rewrite the original thought $r_{\text{orig}}$. The incorrect thoughts are denoted as $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$, based on their origins from $\mathcal{D}_{\text{med}}$, $\mathcal{D}_{\text{easy}}$ and $\mathcal{D}_{\text{hard}}$. These thoughts contain factual errors or invalid reasoning steps, which can mislead $\pi_{\text{base}}$, rendering it unable to solve $x$. Thus, we categorize the properties of all the thoughts we have collected into the following three types: i) $r_{\text{med}}$, $*r_{\text{easy*}}$, and $*r_{\text{hard*}}$: medium-level reasoning processes that are both correct and cognitively suitable for $\pi_{\text{base}}$; ii) $r_{\text{easy}}$ and $r_{\text{hard}}$: easy or hard thoughts that are correct but unsuitable for $\pi_{\text{base}}$; iii) $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$: incorrect reasoning processes with logical flaws or invalid reasoning steps (regardless of the difficulty levels).
@@
Figure 3: An illustration of CogPO, showing the different preference gaps between CoT process pairs and the corresponding mini-tasks.
To define the mini-tasks used for CogPO training, we consider the preference gaps in these three types of thought pairs as follows:
  • Small Gap Mini-task: The thought pairs are from ($*r_{\text{easy*}}$, $r_{\text{easy}}$) and ($*r_{\text{hard*}}$, $r_{\text{hard}}$). Both are correct but differ in complexity (suitable vs. unsuitable for $\pi_{\text{base}}$). We treat $*r_{\text{easy*}}$ and $*r_{\text{hard*}}$ as chosen reasoning processes ($r_{w}$), and $r_{\text{easy}}$ and $r_{\text{hard}}$ as rejected ($r_{l}$).
  • Medium Gap Mini-task: The thought pairs are from ($r_{\text{easy}}$, $\tilde{r}_{\text{easy}}$) and ($r_{\text{hard}}$, $\tilde{r}_{\text{hard}}$). The former are correct but unsuitable, while the latter are completely incorrect. As correctness is more important than suitability for our model, the preference gap of this mini-task should be higher than that in the previous case. For this mini-task, $r_{\text{easy}}$ and $r_{\text{hard}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
  • Large Gap Mini-task: The thought pairs are from ($r_{\text{med}}$, $\tilde{r}_{\text{med}}$), ($*r_{\text{easy*}}$, $\tilde{r}_{\text{easy}}$) and ($*r_{\text{hard*}}$, $\tilde{r}_{\text{hard}}$). Intuitively, the preference gaps should be the largest between suitable and correct thoughts, and incorrect ones. Here, $r_{\text{med}}$, $*r_{\text{easy*}}$ and $*r_{\text{hard*}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
Following our modeling framework, CogPO can be implemented either at the instance level or batch level. For the instance level, each training instance $(x,r_{w},r_{l})$ receives its specific $\beta$ value, as illustrated in Figure 3. The CogPO objective function aggregates these preferences:
$\mathcal{L}_{\text{CogPO}}=-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}}\log\sigma(M_{\beta_{\text{CogPO}}}(x,r_{w},r_{l}))$ (5)
As for the batch level, we require our trainer to sample a mini-task for each batch to optimize. The objective function of CogPO is then rewritten as:
$\mathcal{L}_{\text{CogPO}}=$ $-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{S}}}\log\sigma(M_{\beta_{\text{S}}}(x,r_{w},r_{l}))$
$-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{M}}}\log\sigma(M_{\beta_{\text{M}}}(x,r_{w},r_{l}))$
$-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}_{\text{L}}}\log\sigma(M_{\beta_{\text{L}}}(x,r_{w},r_{l})).$
where $\mathcal{D}_{\text{S}}$, $\mathcal{D}_{\text{M}}$ and $\mathcal{D}_{\text{L}}$ denote the subsets.
Overall, our CogPO algorithm enables granular preference learning: strong regularization ($\beta_{\text{L}}$) for validity discrimination, moderate guidance ($\beta_{\text{M}}$) for suitability alignment, and subtle refinement ($\beta_{\text{S}}$) for reasoning style adaptation. This design provides more control over the thought alignment process, leading to further improvements compared to vanilla SFT distillation or SFT distillation with rewriting based on the CRV system.
Remarks. CogPO can be naturally combined with $\beta$-DPO Wu et al. (2024a). We can redefine the $\beta$ values $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$. Take our instance-level implementation as an example. We define the new instance-level value $\beta_{i}^{*}$ as follows: $\beta_{i}^{*}=\beta_{i}+\alpha\cdot(M_{i}-M_{0})\cdot\beta_{i}$ where $\beta_{i}$ is chosen from $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$ based on the corresponding gap type, $M_{i}$ is the instance-level reward differential, and $M_{0}$ is a predefined threshold as in Wu et al. (2024a). [^7] In our experiment, the combination does not bring substantial improvements, as prior knowledge is more important for our task. Hence, we stick to the usage of $\mathcal{L}_{\text{CogPO}}$.
Dataset/Model Zero-shot SFT CRV+SFT CRV+SFT DPO $\beta$-DPO SimPO CogPO
AIME2024 10.0 20.0 26.7 26.7 23.3 23.3 26.7 30.0
MATH-500 73.6 80.0 83.4 84.0 83.4 83.8 84.2 84.4
GSM8K 89.5 92.3 92.7 91.9 92.6 93.0 92.6 93.3
GPQA Diamond 33.3 37.4 40.9 38.4 40.0 37.4 40.9 40.9
LiveCodeBench V2 30.7 31.3 34.3 34.4 34.4 35.8 36.2 36.6
MMLU 71.9 76.1 76.4 76.5 76.1 76.0 76.5 76.2
OlympiadBench (math-en) 40.1 43.6 45.7 45.8 45.7 46.5 46.0 46.6
@@
Table 1: Performance comparison of various training methods on challenging reasoning benchmarks. The LLM backbone is Qwen2.5-7B-Instruct. The training set is Bespoke-Stratos-17k. Results are shown for zero-shot (without further training), vanilla SFT, CRV (a variant of our approach without thought rewriting), CRV (full implementation), DPO, $\beta$-DPO, SimPO, CogPO methods. DPO, $\beta$-DPO, SimPO and CogPO are conducted on model checkpoints after CRV, using the same preference pair dataset. The metrics are scores of these tasks, with the best results for each dataset in each group marked in bold and the second best underlined.
4 Experiments
To evaluate the effectiveness of the CRV framework and the CogPO algorithm, we conduct a series of experiments on challenging reasoning benchmarks.
4.1 Datasets
In our experiments, we evaluate our work on several benchmarks, including AIME2024888https://maa.org/math-competitions/american-invitational-mathematics-examination-aime, MATH-500 Lightman et al. (2023), GSM8K Cobbe et al. (2021), GPQA Diamond Rein et al. (2023), LiveCodeBench V2 Jain et al. (2024), MMLU Hendrycks et al. (2021), and OlympiadBench (math-en) He et al. (2024). Among these, LiveCodeBench V2 provides a comprehensive evaluation for code generation, MMLU is a large multi-task question-answering benchmark, while the others are mathematical problem sets of varying levels. Since some datasets (e.g., GSM8K) contain their own training sets without deep CoTs, for a unified evaluation, we do not use any training sets from these datasets and instead report the results on the testing sets, regarded as out-of-domain testing. The statistics are shown in the appendix.
For the construction of our training set $*\mathcal{D}_{\text{SFT*}}$, we leverage Bespoke-Stratos-17k [^9] https://huggingface.co/datasets/bespokelabs/Bespoke-Stratos-17k, which contains 17K tuples of questions, reasoning processes, and answers directly distilled from DeepSeek-R1 DeepSeek-AI (2025). We also utilize two recently released CoT datasets to conduct supplementary experiments with details presented in the appendix. The first one is Sky-T1-data-17k [^10] https://github.com/NovaSky-AI/SkyThought, which is distilled from QwQ-32B-Preview, where the reasoning abilities are reported to be weaker than those of Bespoke-Stratos-17k. The second one is OpenThoughts-114k [^11] https://huggingface.co/datasets/open-thoughts/OpenThoughts-114k, which is distilled from DeepSeek-R1 and verified using a data curation recipe. We have chosen not to use some previously released CoT datasets (e.g., OpenLongCoT-SFT [^12] https://huggingface.co/datasets/SimpleBerry/OpenLongCoT-SFT) due to their significantly weaker reasoning abilities, while some of our benchmarks (e.g., AIME2024, OlympiadBench) are extremely challenging, even for human experts.
@@ 4.1 Datasets
Table 2: Comparison between our model and other small reasoning models in the open-source community. Specifically, we train two versions using our approach on Bespoke-Stratos-17k and OpenThoughts-114k, respectively, where the training sets are in line with Bespoke-Stratos-7B and OpenThinker-7B.
Model Backbone (The Critic) AIME2024 MATH-500 GPQA GSM8K LCB OlympiadBench
Qwen2.5-7B-Instruct 13.3 80.2 40.9 92.7 30.5 43.9
Qwen2.5-32B-Instruct 23.3 82.2 39.9 92.9 33.3 45.1
Qwen2.5-72B-Instruct 20.0 81.8 36.4 93.0 30.5 42.0
DeepSeek-R1-Distill-Qwen-32B 26.7 83.4 40.9 92.7 34.3 45.7
@@
Table 3: Comparison using different backbones as the Critic. GPQA: GPQA Diamond. LCB: LiveCodeBench V2. All the results are produced based on the SFT distillation using rewritten training sets without CogPO training.
4.2 Experimental Details
[…]
@@ 4.3 Main Experimental Results and Ablations
Table 1 presents the results of our CRV framework and the CogPO algorithm on various reasoning benchmarks. The training set is Bespoke-Stratos-17k. These results reveal that our CRV framework and the CogPO algorithm significantly enhance the reasoning capabilities of smaller models over traditional methods. In particular, CogPO shows the strongest performance across almost all datasets, indicating its ability to align model thinking processes with their cognitive capacities.
For the ablation studies, we demonstrate that the full implementation of CRV yields gains over its variant CRV, which does not include thought rewriting, in most datasets. This proves the significance of our CRV framework design. However, in a few cases, the performance experiences a slight decline, likely due to the fact that the LLM for the Rethinker inevitably introduces some noise or errors into the training set. Nonetheless, the performance is still higher than vanilla SFT. For alignment training, other algorithms, such as DPO, $\beta$-DPO and SimPO do not necessarily have positive effects on the final scores, illustrating that our CogPO is more suitable training for small reasoning LLMs.

### v2

@@ 3.2.3 Verifier
Finally, we leverage the LLM agent $\pi_{\text{base}}$ to validate the correctness of $r_{\text{med}}$, $r_{\text{easy*}}$, and $r_{\text{hard*}}$ in order to preserve the high quality of the dataset. It predicts whether $\pi_{\text{base}}$ can derive the correct answer $y$ from the rewritten thoughts $r_{\text{easy*}}$ or $r_{\text{hard*}}$. Note that $r_{\text{med}}$ has already been validated as correct in the original dataset, and we send $r_{\text{med}}$ to the Verifier to further ensure data quality.
After verification, incorrect cases are sent back to the Rethinker to be continuously rewritten until they pass verification. In the implementation, cases that fail after three iterations are discarded. For cases that pass verification, we invoke the Critic to make the judgment again (please refer to Figure 2 for the algorithmic flow).
The final SFT dataset is composed of verified medium-level data: $\mathcal{D}_{\text{SFT*}}=\mathcal{D}_{\text{med}}\cup\mathcal{D}_{\text{easy*}}\cup\mathcal{D}_{\text{hard*}}$, where $\mathcal{D}_{\text{med}}$ denotes the verified medium-level data, and $\mathcal{D}_{\text{easy*}}$ and $\mathcal{D}_{\text{hard*}}$ represent the rewritten versions of $\mathcal{D}_{\text{easy}}$ and $\mathcal{D}_{\text{hard}}$ that have passed verification and have been re-rated as medium by the Critic, respectively. $\mathcal{D}_{\text{SFT*}}$ serves as the SFT training set in the CRV stage. Prompt templates used in the CRV system are provided in Appendix C.
@@ 3.3 Cognitive Preference Optimization
The CogPO algorithm aligns CoT processes of smaller LLMs with their inherent cognitive capacities, following the SFT training using the CRV system.
@@ 3.3.1 Preliminaries
Briefly speaking, the CogPO algorithm is extended from DPO Rafailov et al. (2023) and its variants. Let $y_{w}$ and $y_{l}$ be the chosen and rejected responses for an instruction $x$ (not restricted to reasoning problems addressed in this work), respectively. We further denote $\pi_{\theta}$ as the model to be optimized after SFT and $\pi_{\text{ref}}$ as the reference model. DPO seeks to maximize the following margin:
@@
The setting of $\beta$ is critical to the performance of DPO. $\beta$-DPO Wu et al. (2024a) further adjusts $\beta$ according to $M_{\beta}(x,y_{w},y_{l})$, either at the instance level or batch level, allowing the model to adapt $\beta$ based on the reward differential of the input data.
@@ 3.3.2 Algorithmic Description
As noted, DPO and $\beta$-DPO do not require any prior knowledge of how the model learns the user’s preferences. We suggest that this type of prior knowledge is critical for training better small reasoning models, as the cognitive trajectories of large and small models often differ Li et al. (2022); Zhang et al. (2024); Hu et al. (2024), and this may not be directly reflected in the reward differential. We propose CogPO to align reasoning preferences by encoding more prior knowledge and continuously training on a series of mini-tasks.
@@
Figure 3: An illustration of the proposed CogPO algorithm, showing the different preference gaps between CoT pairs and the corresponding mini-tasks.
Dataset/Model Zero-shot SFT CRV+SFT DPO $\beta$-DPO SimPO CogPO
AIME2024 10.0 20.0 26.7 23.3 23.3 26.7 30.0
MATH-500 73.6 80.0 84.0 83.4 83.8 84.2 84.4
GSM8K 89.5 92.3 92.7 92.6 93.0 92.6 93.3
GPQA Diamond 33.3 37.4 40.9 40.0 37.4 40.9 40.9
LiveCodeBench V2 30.7 31.3 34.4 34.4 35.8 36.2 36.6
MMLU 71.9 76.1 76.5 76.1 76.0 76.5 76.5
OlympiadBench (math-en) 40.1 43.6 45.8 45.7 46.5 46.0 46.6
@@
Table 2: Performance comparison of various training methods. The LLM backbone is Qwen2.5-7B-Instruct, and the training set is Bespoke-Stratos-17k. Results are shown for zero-shot (without further training), SFT, CRV+SFT, DPO, $\beta$-DPO, SimPO, and CogPO. DPO, $\beta$-DPO, SimPO, and CogPO are conducted on the same model checkpoints of CRV+SFT, using the same preference pair dataset. The metrics represent scores for these tasks, with the best results for each dataset in each group marked in bold and the second-best underlined.
@@
Table 3: Comparison between our model and other small reasoning models in the open-source community. Specifically, we train two versions using our approach on Bespoke-Stratos-17k and OpenThoughts-114k, respectively, where the two training sets are the same as Bespoke-Stratos-7B and OpenThinker-7B, respectively.
We leverage the Rethinker in CRV to also generate incorrect reasoning processes when prompted to rewrite the original thought $r_{\text{orig}}$ (prompt template is provided in Appendix C). The incorrect thoughts are denoted as $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$, and $\tilde{r}_{\text{hard}}$, based on their origin from $\mathcal{D}_{\text{med}}$, $\mathcal{D}_{\text{easy}}$, and $\mathcal{D}_{\text{hard}}$. These thoughts contain factual errors or invalid reasoning steps, which can mislead $\pi_{\text{base}}$, rendering it impossible to solve $x$. Thus, we categorize the properties of all the thoughts collected into the following three types: i) $r_{\text{med}}$, $r_{\text{easy*}}$, and $r_{\text{hard*}}$: medium-level reasoning processes that are both correct and cognitively suitable for $\pi_{\text{base}}$; ii) $r_{\text{easy}}$ and $r_{\text{hard}}$: easy or hard thoughts that are correct but unsuitable for $\pi_{\text{base}}$; iii) $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$, and $\tilde{r}_{\text{hard}}$: incorrect reasoning processes with logical flaws or invalid reasoning steps (regardless of the difficulty levels). To define the mini-tasks used for CogPO training, we consider the preference gaps in these three types of CoT pairs as follows:
  • Small Gap Mini-task: The pairs are ($r_{\text{easy*}}$, $r_{\text{easy}}$) and ($r_{\text{hard*}}$, $r_{\text{hard}}$). Both are correct but differ in complexity (suitable vs. unsuitable for $\pi_{\text{base}}$). We treat $r_{\text{easy*}}$ and $r_{\text{hard*}}$ as chosen reasoning processes ($r_{w}$), and $r_{\text{easy}}$ and $r_{\text{hard}}$ as rejected ($r_{l}$).
  • Medium Gap Mini-task: The pairs are ($r_{\text{easy}}$, $\tilde{r}_{\text{easy}}$) and ($r_{\text{hard}}$, $\tilde{r}_{\text{hard}}$). The former are correct but unsuitable, while the latter are completely incorrect. As correctness is more important than suitability for our model, the preference gap of this mini-task should be higher than that in the previous case. For this mini-task, $r_{\text{easy}}$ and $r_{\text{hard}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{easy}}$ and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
  • Large Gap Mini-task: The pairs are ($r_{\text{med}}$, $\tilde{r}_{\text{med}}$), ($r_{\text{easy*}}$, $\tilde{r}_{\text{easy}}$), and ($r_{\text{hard*}}$, $\tilde{r}_{\text{hard}}$). Intuitively, the preference gaps should be the largest between suitable and correct thoughts and incorrect ones. Here, $r_{\text{med}}$, $r_{\text{easy*}}$, and $r_{\text{hard*}}$ are treated as $r_{w}$, while $\tilde{r}_{\text{med}}$, $\tilde{r}_{\text{easy}}$, and $\tilde{r}_{\text{hard}}$ are treated as $r_{l}$.
Following our modeling framework, each training instance $(x,r_{w},r_{l})$ receives its specific $\beta$ value, as illustrated in Figure 3. The CogPO objective function aggregates these preferences:
$\mathcal{L}_{\text{CogPO}}=-\mathbb{E}_{(x,r_{w},r_{l})\sim\mathcal{D}}\log\sigma(M_{\beta_{\text{CogPO}}}(x,r_{w},r_{l})),$ (3)
Overall, our CogPO algorithm enables granular preference learning: strong regularization ($\beta_{\text{L}}$) for validity discrimination, moderate guidance ($\beta_{\text{M}}$) for suitability alignment, and subtle refinement ($\beta_{\text{S}}$) for reasoning style adaptation. This design provides more control over the alignment process, leading to further improvements on the basis of SFT (using the CRV system).
Remarks. CogPO can be naturally combined with $\beta$-DPO Wu et al. (2024a). We can redefine the $\beta$ values $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$ as follows:
$\beta_{i}^{*}=\beta_{i}+\alpha\cdot(M_{i}-M_{0})\cdot\beta_{i}$ (4)
where $\beta_{i}$ is chosen from $\{\beta_{\text{S}},\beta_{\text{M}},\beta_{\text{L}}\}$ based on the corresponding gap type, $M_{i}$ is the instance-level reward differential, and $M_{0}$ is a predefined threshold as in Wu et al. (2024a). [^7] In our experiments, this combination does not yield substantial improvements, as prior knowledge is more important for our task. Hence, we stick to using $\mathcal{L}_{\text{CogPO}}$.
4 Experiments
To evaluate the effectiveness of the CRV framework and the CogPO algorithm, we conduct a series of experiments on several challenging reasoning benchmarks. Due to space limitations, the datasets and experimental settings are described in Appendix A.1 and A.2.
Model Backbone (The Critic) AIME2024 MATH-500 GPQA-D GSM8K LCB V2 OlympiadBench
Qwen2.5-7B-Instruct 13.3 80.2 40.9 92.3 30.5 43.9
Qwen2.5-32B-Instruct 23.3 82.2 39.9 92.6 33.3 45.1
Qwen2.5-72B-Instruct 20.0 81.8 36.4 92.7 30.5 42.0
DeepSeek-R1-Distill-Qwen-32B 26.7 84.0 40.9 92.7 34.4 45.8
@@ 4 Experiments
Table 4: Comparison using different backbones as the Critic. All results are produced using CRV+SFT without CogPO on Bespoke-Stratos-17k.
@@ 4.1 Main Experimental Results and Ablations
We choose Bespoke-Stratos-17k as our training set. Table 2 presents the results of our CRV framework and the CogPO algorithm on various reasoning benchmarks. CRV+SFT outperforms direct SFT on all benchmarks. Building on CRV+SFT, CogPO further enhances the model’s reasoning ability, surpasses other preference‐optimization algorithms, and ultimately achieves the strongest performance, demonstrating its effectiveness in aligning the model’s reasoning processes with its cognitive capacities. These results reveal that our CRV+CogPO framework effectively enhances the reasoning capabilities of smaller models, outperforming other methods by a large margin.